# Oryx MLLM: On-Demand Spatial-Temporal Understanding at Arbitrary Resolution

### v1

@@ 4.3 Long-Form Temporal Understanding
Model Size MLVU LongVideoBench VideoMME-Long w/o subs w subs Proprietary Models GPT-4V (OpenAI, 2023b) - 49.2 60.7 53.5 56.9 GPT-4o (OpenAI, 2024) - 64.6 66.7 65.3 72.1 Gemini-1.5-Pro (GeminiTeam, 2024) - - 64.4 67.4 77.4 Open-Sourced Video MLLMs VideoLLaMA2 (Cheng et al., 2024) 7B 48.5 - 42.1 43.8 LongVA (Zhang et al., 2024a) 7B 56.3 - 46.2 47.6 LLaVA-OneVision (Li et al., 2024a) 7B 64.7 - - - Kangaroo (Liu et al., 2024e) 8B 61.0 54.8 46.6 49.3 LongVILA (Xue et al., 2024) 8B - - 39.7 - VideoCCAM (Fei et al., 2024) 14B 63.1 - 46.7 49.9 LLaVA-Next-Video (Zhang et al., 2024c) 34B - 50.5 - - PLLaVA (Xu et al., 2024a) 34B - 53.2 - - VILA-1.5 (Lin et al., 2023b) 40B 56.7 - 53.8 55.7 LLaVA-OneVision (Li et al., 2024a) 72B 66.4 61.3 60.0 62.4 Oryx 7B 67.5 55.3 50.3 55.8 Oryx 34B 70.8 62.2 53.9 58.0
@@ 4.3.1 Long-Form Video Benchmarks
Setup. We select three mainstream and representative benchmarks specifically designed for long video understanding, ensuring a comprehensive evaluation of long video inputs. MLVU (Zhou et al., 2024), encompasses videos ranging from 3 minutes to 2 hours and includes 9 distinct tasks that assess both global and local information within the video content. LongVideoBench (Wu et al., 2024) presents a primary challenge of retrieving and reasoning over a dataset comprising 3k long video inputs. Additionally, we utilize the long video subset of the VideoMME (Fu et al., 2024) benchmark, which features videos with lengths ranging from 30 minutes to 60 minutes.
Results. Results are shown in Table 2, which highlights the efficacy of our unified and on-demand design across varying temporal lengths and our further efforts in the context of long video retrieval. The Oryx model exhibits a remarkable capability in understanding long-form video content. Specifically, our Oryx-7B model surpasses all existing 7B model series on long video benchmarks. Furthermore, the Oryx-34B model showcases strong performance across larger MLLMs, achieving a mean accuracy improvement of 4.4% and 0.9% over previous state-of-the-art models equipped with 72B parameter LLMs on the MLVU (Zhou et al., 2024) and LongVideoBench (Wu et al., 2024) benchmarks, respectively. Notably, the Oryx-34B model also outperforms GPT-4o (OpenAI, 2024) on the challenging MLVU (Zhou et al., 2024) benchmark by a margin of 6.2%, underscoring its advanced capabilities in long video understanding.
@@ 4.3.2 Video Needle-In-A-Haystack
Model Size MMBench MMMU DocVQA OCRBench AI2D TextVQA Deepseek-VL (Lu et al., 2024) 7B 73.2 36.6 - 456 - 64.7 Monkey (Li et al., 2024d) 7B 72.4 40.7 - 534 68.5 - LLaVA-NeXT (Liu et al., 2024c) 8B 72.1 41.7 78.2 531 71.6 - Bunny-LLama3 (He et al., 2024) 8B 77.2 43.3 - 444 69.4 - Cambrian-1 (Tong et al., 2024) 8B 75.9 42.7 77.8 624 73.6 71.7 VILA-1.5 (Lin et al., 2023b) 8B 75.3 38.6 - - - 68.5 Idefics2 (Laurençon et al., 2024) 8B 76.7 43.0 - - - 73.0 Yi-VL (Young et al., 2024) 34B - 45.1 - 290 65.9 - LLaVA-NeXT (Liu et al., 2024c) 34B 79.3 49.7 84.0 574 74.9 - Cambrian-1 (Tong et al., 2024) 34B 81.4 49.7 75.5 600 79.7 76.7 VILA-1.5 (Lin et al., 2023b) 40B 82.4 51.9 - - - 73.4 Oryx 7B 81.4 43.9 89.0 672 78.5 75.0 Oryx 34B 84.5 50.3 91.4 743 81.0 77.8
@@
Image Benchmarks. We select a diverse set of mainstream and representative image benchmarks to evaluate the model’s proficiency in image understanding. Specifically, we included MMBench (Liu et al., 2023a) and MMMU (Yue et al., 2024) to assess general image understanding capabilities, and DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2023b), AI2D (Kembhavi et al., 2016), and TextVQA (Singh et al., 2019) to evaluate the model’s performance on specific tasks such as document recognition, OCR, text understanding tasks, etc. The results are summarized in Table 3. Notably, the Oryx model maintains pioneering results on image benchmarks, such as an 84.5% mean accuracy on MMBench (Liu et al., 2023a) and a 91.4% accuracy on DocVQA (Mathew et al., 2021). Such results demonstrate the effectiveness of our method in comprehending images with more simple and lightweight training pipelines, data curation, and strategies compared with concurrent works.
@@
Model Size   METEOR   ROUHE-L   CIDEr   BLEU-1   BLEU-2 3D-Specific Models VoteNet (Qi et al., 2019) - 11.4 29.8 54.7 28.0 16.7 ScanQA (Azuma et al., 2022) - 11.5 30 55.4 26.9 16.6 ScanRefer (Chen et al., 2020) - 13.1 33.3 64.9 30.2 20.4 3D-LLM (Hong et al., 2023) - 14.5 35.7 69.4 39.3 25.2 General Open-Source MLLMs BLIP2 (Li et al., 2023) - 11.3 26.6 45.7 29.7 16.2 Flamingo (Alayrac et al., 2022) 7B 11.3 31.1 55 25.6 15.2 Mantis (Jiang et al., 2024) 7B - 16.1 - - - LLaVA-Next-Interleave (Li et al., 2024b) 14B - 34.5 - - - LLaVA-OneVision (Li et al., 2024a) 72B - 35.8 - - - Oryx 7B 14.0 35.1 68.7 35.3 22.3 Oryx 34B 15.0 37.3 72.3 38.0 24.6
@@ Appendix B Training Details
Stage 2. For stage 2, we continuously train the Oryx model from the multi-modal LLMs in stage 1. We randomly sample around 600k image data from the supervised fine-tuning stage in stage 1 and add additional 650k temporal and 3D data from open-source multi-modal datasets, resulting in an overall number of 1.2M further supervised fine-tuning data. In the more general stage, we increase the restriction for image pixels to $1536^{2}$ to meet the longer sequential length in temporal data. We maintain the aspect ratio of video data while normalizing each frame to the minimum size of $288^{2}$ pixels and the maximum size of $480^{2}$ pixels, therefore the token length before compression module ranges from 324 to 900. We adopt $1\times 1$ path for the image data, $2\times 2$ pooling path for the multi-frame data including video and 3D-relevant data, and $4\times 4$ pooling path for the extremely long video needle-in-the-haystack retrieval data. We maintain most of the training hyper-parameters identical to stage 1, with a total batch size of 128, a learning rate of 2e-5 for Oryx-7B, and a learning rate of 1e-5 for Oryx-34B. We sample 1 frame per second for video data and set the max frame number at 64 frames. We uniformly sample the frames among all the frames if the number exceeds the upper bound. The maximum sequence length is set to 16384.

### v2

@@ 4.3 Long-Form Temporal Understanding
Model Size MLVU LongVideoBench VideoMME-Long w/o subs w subs Proprietary Models GPT-4V (OpenAI, 2023b) - 49.2 60.7 53.5 56.9 GPT-4o (OpenAI, 2024) - 64.6 66.7 65.3 72.1 Gemini-1.5-Pro (GeminiTeam, 2024) - - 64.4 67.4 77.4 Open-Sourced Video MLLMs VideoLLaMA2 (Cheng et al., 2024) 7B 48.5 - 42.1 43.8 LongVA (Zhang et al., 2024a) 7B 56.3 - 46.2 47.6 LLaVA-OneVision (Li et al., 2024a) 7B 64.7 - - - Kangaroo (Liu et al., 2024e) 8B 61.0 54.8 46.6 49.3 LongVILA (Xue et al., 2024) 8B - - 39.7 - VideoCCAM (Fei et al., 2024) 14B 63.1 - 46.7 49.9 LLaVA-Next-Video (Zhang et al., 2024c) 34B - 50.5 - - PLLaVA (Xu et al., 2024a) 34B - 53.2 - - VILA-1.5 (Lin et al., 2023b) 40B 56.7 - 53.8 55.7 LLaVA-OneVision (Li et al., 2024a) 72B 66.4 61.3 60.0 62.4 Oryx 7B 67.5 55.3 50.3 55.8 Oryx 34B 70.8 62.2 53.9 58.0 Oryx-1.5 7B 67.5 56.3 51.2 58.3 Oryx-1.5 32B 72.3 62.0 59.1 69.7
@@ 4.3.1 Long-Form Video Benchmarks
Setup. We select three mainstream and representative benchmarks specifically designed for long video understanding, ensuring a comprehensive evaluation of long video inputs. MLVU (Zhou et al., 2024), encompasses videos ranging from 3 minutes to 2 hours and includes 9 distinct tasks that assess both global and local information within the video content. LongVideoBench (Wu et al., 2024) presents a primary challenge of retrieving and reasoning over a dataset comprising 3k long video inputs. Additionally, we utilize the long video subset of the VideoMME (Fu et al., 2024a) benchmark, which features videos with lengths ranging from 30 minutes to 60 minutes.
Results. Results are shown in Table 2, which highlights the efficacy of our unified and on-demand design across varying temporal lengths and our further efforts in the context of long video retrieval. The Oryx model exhibits a remarkable capability in understanding long-form video content. Specifically, our Oryx-1.5-7B model surpasses all existing 7B model series on long video benchmarks. Furthermore, the Oryx-1.5-32B model showcases strong performance across larger MLLMs, achieving a mean accuracy improvement of 5.9% and 0.7% over previous state-of-the-art models equipped with 72B parameter LLMs on the MLVU (Zhou et al., 2024) and LongVideoBench (Wu et al., 2024) benchmarks, respectively. Notably, the Oryx-1.5-32B model also outperforms GPT-4o (OpenAI, 2024) on the challenging MLVU (Zhou et al., 2024) benchmark by a margin of 7.7%, underscoring its advanced capabilities in long video understanding.
@@ 4.3.2 Video Needle-In-A-Haystack
Model Size MMBench MMMU DocVQA OCRBench AI2D TextVQA Deepseek-VL (Lu et al., 2024) 7B 73.2 36.6 - 456 - 64.7 Monkey (Li et al., 2024d) 7B 72.4 40.7 - 534 68.5 - LLaVA-NeXT (Liu et al., 2024c) 8B 72.1 41.7 78.2 531 71.6 - Bunny-LLama3 (He et al., 2024) 8B 77.2 43.3 - 444 69.4 - Cambrian-1 (Tong et al., 2024) 8B 75.9 42.7 77.8 624 73.6 71.7 VILA-1.5 (Lin et al., 2023b) 8B 75.3 38.6 - - - 68.5 Idefics2 (Laurençon et al., 2024) 8B 76.7 43.0 - - - 73.0 Yi-VL (Young et al., 2024) 34B - 45.1 - 290 65.9 - LLaVA-NeXT (Liu et al., 2024c) 34B 79.3 49.7 84.0 574 74.9 - Cambrian-1 (Tong et al., 2024) 34B 81.4 49.7 75.5 600 79.7 76.7 VILA-1.5 (Lin et al., 2023b) 40B 82.4 51.9 - - - 73.4 VITA (Fu et al., 2024b) 8$\times$7B 71.8 47.3 - 678 73.1 - LLaVA-OneVision (Li et al., 2024a) 72B 85.6 56.8 93.1 - 85.6 - Oryx 7B 81.4 43.9 89.0 672 78.5 75.0 Oryx 34B 84.5 50.3 91.4 743 81.0 77.8 Oryx-1.5 7B 81.3 47.1 90.1 713 79.7 75.7 Oryx-1.5 32B 86.3 56.1 92.7 746 83.2 78.3
@@
Image Benchmarks. We select a diverse set of mainstream and representative image benchmarks to evaluate the model’s proficiency in image understanding. Specifically, we included MMBench (Liu et al., 2023a) and MMMU (Yue et al., 2024) to assess general image understanding capabilities, and DocVQA (Mathew et al., 2021), OCRBench (Liu et al., 2023b), AI2D (Kembhavi et al., 2016), and TextVQA (Singh et al., 2019) to evaluate the model’s performance on specific tasks such as document recognition, OCR, text understanding tasks, etc. The results are summarized in Table 3. Notably, the Oryx model maintains pioneering results on image benchmarks, such as an 86.3% mean accuracy on MMBench (Liu et al., 2023a) and a 92.7% accuracy on DocVQA (Mathew et al., 2021). Such results demonstrate the effectiveness of our method in comprehending images with more simple and lightweight training pipelines, data curation, and strategies compared with concurrent works.
@@
Model Size   METEOR   ROUGE-L   CIDEr   BLEU-1   BLEU-2 3D-Specific Models VoteNet+MCAN (Qi et al., 2019) - 11.4 29.8 54.7 28.0 16.7 ScanQA (Azuma et al., 2022) - 11.5 30 55.4 26.9 16.6 ScanRefer+MCAN (Chen et al., 2020) - 13.1 33.3 64.9 30.2 20.4 3D-LLM (Hong et al., 2023) - 14.5 35.7 69.4 39.3 25.2 General Open-Source MLLMs BLIP2 (Li et al., 2023) - 11.3 26.6 45.7 29.7 16.2 Flamingo (Alayrac et al., 2022) 7B 11.3 31.1 55 25.6 15.2 Mantis (Jiang et al., 2024) 7B - 16.1 - - - LLaVA-Next-Interleave (Li et al., 2024b) 14B - 34.5 - - - LLaVA-OneVision (Li et al., 2024a) 72B - 35.8 - - - Oryx 7B 14.5 35.5 69.1 35.8 24.4 Oryx 34B 15.0 37.3 72.3 39.6 26.7 Oryx-1.5 7B 15.2 38.4 73.5 38.6 24.6 Oryx-1.5 32B 15.3 38.4 74.3 38.8 24.4
@@ Appendix B Training Details
Stage 2. For stage 2, we continuously train the Oryx model from the multi-modal LLMs in stage 1. We randomly sample around 600k image data from the supervised fine-tuning stage in stage 1 and add an additional 650k temporal and 3D data from open-source multi-modal datasets, resulting in an overall number of 1.2M further supervised fine-tuning data. In the more general stage, we increase the restriction for image pixels to $1536^{2}$ to meet the longer sequential length in temporal data. We maintain the aspect ratio of video data while normalizing each frame to the minimum size of $288^{2}$ pixels and the maximum size of $480^{2}$ pixels, therefore the token length before the compression module ranges from 324 to 900. We adopt $1\times 1$ path for the image data, $2\times 2$ pooling path for the multi-frame data including video and 3D-relevant data, and $4\times 4$ pooling path for the extremely long video needle-in-the-haystack retrieval data. We maintain most of the training hyper-parameters identical to stage 1, with a total batch size of 128, a learning rate of 2e-5 for Oryx-7B, and a learning rate of 1e-5 for Oryx-34B and Oryx-1.5-32B. We sample 1 frame per second for video data and set the max frame number to 64 frames. We uniformly sample the frames among all the frames if the number exceeds the upper bound. The maximum sequence length is set to 16384.